OpenStudy (ffrocks3):

Find the equation in standard form which passes through (-2, -7) and (1, -1). You must show all work.

OpenStudy (texaschic101):

first we find the slope by using the slope formula : slope (m) = (y2 - y1) / (x2 - x1) (-2,-7)....x1 = -2 and y1 = -7 (1,-1)...x2 = 1 and y2 = -1 now we sub slope(m) = (-1 - (-7) / (1 - (-2) slope(m) = (-1 + 7) / (1 + 2) slope(m) = 6/3 which reduces to 2 now we use y = mx + b.....and since we know the slope(m) = 2, we now need to find the y intercept(b) slope(m) = 2 you can use either of your sets of points...(-2,-7)...x = -2 and y = -7 now we sub y = mx + b -7 = 2(-2) + b -7 = -4 + b -7 + 4 = b -3 = b so your equation is : y = 2x - 3 but we need it in standard form Ax + By = C y = 2x - 3 --- subtract 2x from both sides -2x + y = -3 --- multiply by -1 to make x positive 2x - y = 3 <== standard form

OpenStudy (ffrocks3):

Ty So Much! That Really Helped Me. :)

OpenStudy (texaschic101):

good to hear...glad I could help :)